# HAMLET: Hyperadaptive Agent-based Modeling for Live Embodied Theatrics

### v1

@@ F.2 Three Responding Strategies: Description and Definition
Slow (Considered Deliberation)
@@
Silence (Strategic Non-Response)
@@ F.3 Annotation Process and Example
Table 7: An example of labeling the PAD model’s strategic choice. The annotator’s task is to determine the most fitting strategy, which in this case is clearly a Slow Response.
@@ F.4 Important Considerations
Appendix G G. Training and Evaluation Details for PAD and HAMLETJudge
@@ G.1 Training Settings and Hyperparameters
Both PAD and HAMLETJudge models were fine-tuned based on Qwen3-8B using the LLaMA-Factory framework. Training was conducted with ZeRO-3 across 8 NVIDIA A800-SXM4-80GB GPUs, utilizing bfloat16 precision. The maximum sequence length was set to 65,536, and the total batch size was 128.
For PAD, we performed supervised fine-tuning (SFT) on the training set for 3 epochs. The optimizer was configured with a learning rate of 1e-5, a cosine learning rate scheduler, and a warmup ratio of 0.1.
For HAMLETJudge, we adopted a two-stage training paradigm. In the first stage, we conducted supervised fine-tuning on the first half of the training set, using the same optimization setup as PAD (learning rate of 1e-5, cosine scheduler, and 0.1 warmup ratio). In the second stage, leveraging the pairwise nature of the dataset, we restructured the remaining half of the training set into chosen–rejected format and applied Direct Preference Optimization (DPO). The DPO stage was trained for 2 epochs with a learning rate of 5e-6, a KL regularization coefficient $\beta=0.1$, and the same cosine learning rate scheduler.
@@ G.2 Data Statistics and Holdout Method
The complete PAD dataset comprises two main categories: Single Tool Use, which includes Fast, Slow, and Silence responses, and Multi Tool Use, which involves combinations such as Action + Fast, Action + Slow, and Action + Silence. Detailed statistics for each category are presented in Table 8.
PAD Single Tool Use Multi Tool Use
Fast Slow Silence Action+Fast Action+Slow Action+Silence
@@
Mapping Rule speech [thinking]+speech no speech (action)+speech (action)+[thinking]+speech (action)
Table 8: Statistics of the PAD Dataset.
The Mapping Rule row in Table 8 illustrates the expected response formats for AI actors under each tool-use strategy. The core objective of PAD is to train agents in selecting appropriate response strategies, namely Fast, Slow, or Silence. In multi-tool use settings, the Action Tool is further annotated with a verb-object pair to reflect its parametrized usage in combination with a primary response strategy.
@@
Given these characteristics, we consider the Action Tool as an auxiliary enhancement to the core response strategy in PAD, rather than a primary focus of quantitative evaluation. The dataset was partitioned via the holdout method (95% training, 5% testing), with stratification to preserve class distribution.
On the other hand, the HAMLETJudge dataset is designed to evaluate drama performance across three key dimensions: CP (Character Performance), NQ (Narrative Quality), and IE (Interaction Experience). For each LLM pair—e.g., Claude-3.5-sonnet vs. GPT-4.5-preview—75 pairwise comparisons are conducted per dimension in both Chinese and English, resulting in 150 annotated instances per dimension for each model pair.
We select 7 representative LLMs exhibiting diverse levels of drama performance, as demonstrated on the HAMLET Leaderboard: Claude-4-sonnet-Thinking, DeepSeek-R1-0528, Qwen3-235B-A22B, Gemini-2.5-pro, GPT-4o, Qwen3-8B, and LLaMA-3.1-8B. This leads to a total of $\binom{7}{2}=21$ unique model pairs. With 150 data points per pair and 3 evaluation dimensions, the final dataset comprises $21\times 150\times 3=9{,}450$ annotated instances.
@@
HAMLETJudge #CP #NQ #IE #Total
@@
Table 9: Statistics of the HAMLETJudge Dataset.
@@
G.3 PAD Latency Penalty
As discussed in the PAD reliability experiments, although models that output reasoning tokens before invoking tool calls generally achieve better performance, this strategy may introduce intolerable extra latency in real-time drama settings. To address this trade-off, we introduce a latency penalty to balance PAD evaluation results.
@@ G.3 PAD Latency Penalty
PAD 0.36 0
Table 10: Average Latency per Evaluation Instance and Corresponding Penalty. Note: Penalty is mapped based on latency: 0 for $<$2s, 0.05 for 2–10s, 0.10 for 10–30s, and 0.15 for $\geq$30s.
@@
Appendix H H. Prompts Used in HAMLET
This section shows the prompt design in HAMLET multi-agent workflow. The content between { } is variants or reference templates. Responsibilities of all agents and their corresponding prompt designs are as follows:
H.1. Agents in Offline Planning
@@
H.2. Agents in Online Performance
During the online performance, agents work together to create a dynamic and coherent drama. The Critic (Table 17) evaluates the performance quality by comparing different scene generations against specific criteria. The Narrator ensures all physical interactions are logical and realistic, updating the environment’s state based on the truth. The Planner (Table 18) lays out the narrative paths between key plot points and makes sure actors follow a coherent sequence, preventing them from skipping ahead in the story. The Transfer agent (Table 19) monitors the dialogue to identify when a specific condition, which is called a flag, for advancing the plot has been met. If the story stalls, the Advancer (Table 20) steps in, giving actors direct instructions to move the plot forward. Finally, the Actor agents (Table 21 and Table 22) bring the characters to life. They first use the PAD module to assess the situation and select a response strategy. Then, they generate the character’s performance, which can be a mix of dialogue, actions, and internal thoughts.
@@ H.2. Agents in Online Performance
Prompt: Actor (PAD Strategy) You are a role-playing expert that can perceive the surrounding environment and decide the appropriate responding strategy to the current speaker under considerable and comprehensive consideration. You may call one or more functions to assist with the user query. You are provided with function signatures within <tools></tools> XML tags: <tools>{pad_tools_str}</tools> For each function call, return a json object with function name and arguments within <tool_call></tool_call> XML tags:<tool_call>{{"name": <function-name>, "arguments": <argsjsonobject>}}</tool_call> You are now in {environment_description}, current actors are {actor_list}, dialogue history is {current_chat_history}, interactable objects are {props}, your persona, memory, goal and relationships are described in {profile}. Now Current Speaker says: {last_sentence}. Please decide how to respond with most appropriate tool use or tool use combination.
Table 21: The prompt for the Actor’s PAD (Perceive, Analyze, Decide) module.
@@
Appendix I I. Author Contributions
Sizhou Chen and Shufan Jiang contributed equally to this work. Sizhou Chen and Shufan Jiang jointly conceived the study, designed the method, and wrote the manuscript. Sizhou Chen prepared the figures. Shufan Jiang performed the large-scale model experiments. All other authors contributed to the manuscript review. All authors have read and agreed to the published version of the manuscript.

### v2

@@ F.2.1 Fast (Intuitive Reaction)
F.2.2 Slow (Considered Deliberation)
@@
F.2.3 Silence (Strategic Non-Response)
@@ F.3 Annotation Process and Example
Table 7: An example of labeling the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: model’s strategic choice. The annotator’s task is to determine the most fitting strategy, which in this case is clearly a Slow Response.
@@ F.4 Important Considerations
Appendix G Training and Evaluation Details for 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
@@ G.1 Training Settings and Hyperparameters
Both 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: models were fine-tuned based on Qwen3-8B using the LLaMA-Factory framework. Training was conducted with ZeRO-3 across 8 NVIDIA A800-SXM4-80GB GPUs, utilizing bfloat16 precision. The maximum sequence length was set to 65,536, and the total batch size was 128.
For 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, we performed supervised fine-tuning (SFT) on the training set for 3 epochs. The optimizer was configured with a learning rate of 1e-5, a cosine learning rate scheduler, and a warmup ratio of 0.1.
For 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, we adopted a two-stage training paradigm. In the first stage, we conducted supervised fine-tuning on the first half of the training set, using the same optimization setup as 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: (learning rate of 1e-5, cosine scheduler, and 0.1 warmup ratio). In the second stage, leveraging the pairwise nature of the dataset, we restructured the remaining half of the training set into chosen–rejected format and applied Direct Preference Optimization (DPO). The DPO stage was trained for 2 epochs with a learning rate of 5e-6, a KL regularization coefficient $\beta=0.1$, and the same cosine learning rate scheduler.
@@ G.2 Data Statistics and Holdout Method
The complete 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset comprises two main categories: Single Tool Use, which includes Fast, Slow, and Silence responses, and Multi Tool Use, which involves combinations such as Action + Fast, Action + Slow, and Action + Silence. Detailed statistics for each category are presented in Table 8.
0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Single Tool Use Multi Tool Use
Fast Slow Silence
Action+Fast
Action+Slow
Action+Silence
@@
Mapping Rule speech
[thinking]+speech
no speech
(action)+speech
(action)+
[thinking]+speech
Table 8: Statistics of the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Dataset.
The Mapping Rule row in Table 8 illustrates the expected response formats for AI actors under each tool-use strategy. The core objective of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is to train agents in selecting appropriate response strategies, namely Fast, Slow, or Silence. In multi-tool use settings, the Action Tool is further annotated with a verb-object pair to reflect its parametrized usage in combination with a primary response strategy.
@@
Given these characteristics, we consider the Action Tool as an auxiliary enhancement to the core response strategy in 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, rather than a primary focus of quantitative evaluation. The dataset was partitioned via the holdout method (95% training, 5% testing), with stratification to preserve class distribution.
On the other hand, the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset is designed to evaluate drama performance across three key dimensions: CP (Character Performance), NQ (Narrative Quality), and IE (Interaction Experience). For each LLM pair—e.g., Claude-3.5-sonnet vs. GPT-4.5-preview—75 pairwise comparisons are conducted per dimension in both Chinese and English, resulting in 150 annotated instances per dimension for each model pair.
We select 7 representative LLMs exhibiting diverse levels of drama performance, as demonstrated on the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Leaderboard: Claude-4-sonnet-Thinking, DeepSeek-R1-0528, Qwen3-235B-A22B, Gemini-2.5-pro, GPT-4o, Qwen3-8B, and LLaMA-3.1-8B. This leads to a total of $\binom{7}{2}=21$ unique model pairs. With 150 data points per pair and 3 evaluation dimensions, the final dataset comprises $21\times 150\times 3=9{,}450$ annotated instances.
@@
0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: #CP #NQ #IE Total
@@
Table 9: Statistics of the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Dataset.
@@
G.3 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Latency Penalty
As discussed in the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: reliability experiments, although models that output reasoning tokens before invoking tool calls generally achieve better performance, this strategy may introduce intolerable extra latency in real-time drama settings. To address this trade-off, we introduce a latency penalty to balance 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: evaluation results.
@@
0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: 0.36 0
Table 10: Average Latency per Evaluation Instance and Corresponding Penalty. Note: Penalty is mapped based on latency: 0 for <<2s, 0.05 for 2–10s, 0.10 for 10–30s, and 0.15 for $\geq$30s.
@@
Appendix H Prompts Used in 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
This section shows the prompt design used in 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: multi-agent workflow. The content between { } is variants or reference templates. Responsibilities of all agents and their corresponding prompt designs are as follows:
H.1 Agents in Offline Planning
@@
H.2 Agents in Online Performance
During the online performance, agents work together to create a dynamic and coherent drama. The Critic (Table 17) evaluates the performance quality by comparing different scene generations against specific criteria. The Narrator ensures all physical interactions are logical and realistic, updating the environment’s state based on the truth. The Planner (Table 18) lays out the narrative paths between key plot points and makes sure actors follow a coherent sequence, preventing them from skipping ahead in the story. The Transfer agent (Table 19) monitors the dialogue to identify when a specific condition, which is called a flag, for advancing the plot has been met. If the story stalls, the Advancer (Table 20) steps in, giving actors direct instructions to move the plot forward. Finally, the Actor agents (Table 21 and Table 22) bring the characters to life. They first use the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: module to assess the situation and select a response strategy. Then, they generate the character’s performance, which can be a mix of dialogue, actions, and internal thoughts.
Appendix I Large Language Models Usage Statement
During the preparation of this work, we utilized Large Language Models (LLMs), such as ChatGPT, as a writing-assistant tool. The use of LLMs was exclusively limited to improving the grammar, clarity, and readability of the text. All the scientific ideas, experimental designs, results, and conclusions presented in this paper were conceived and formulated entirely by the human authors. The authors take full responsibility for the content of this paper.
@@ Appendix I Large Language Models Usage Statement
Prompt: PAD module) You are a role-playing expert that can perceive the surrounding environment and decide the appropriate responding strategy to the current speaker under considerable and comprehensive consideration. You may call one or more functions to assist with the user query. You are provided with function signatures within <tools></tools> XML tags: <tools>{pad_tools_str}</tools> For each function call, return a json object with function name and arguments within <tool_call></tool_call> XML tags:<tool_call>{{"name": <function-name>, "arguments": <argsjsonobject>}}</tool_call> You are now in {environment_description}, current actors are {actor_list}, dialogue history is {current_chat_history}, interactable objects are {props}, your persona, memory, goal and relationships are described in {profile}. Now Current Speaker says: {last_sentence}. Please decide how to respond with most appropriate tool use or tool use combination.
Table 21: The prompt for the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: module.